OpenStudy (anonymous):

The table shows the annual consumption of cheese per person in the United States for selected years in the 20th century. YEAR: POUNDS CONSUMED: 1908 5.027 1923 8.768 1960 12.29 1981 15.485 ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ y = –0.0000541x^3 + 0.00789x^2 – 0.452x – 1.888 y = –0.0000541x^3 – 0.00789x^2 – 0.452x + 1.888 y = 0.00789x^3 – 0.0000541x^2 + 1.888x + 0.452 y = 0.0000541x^3 – 0.00789x^2 + 0.452x + 1.888

OpenStudy (anonymous):

and the question?

OpenStudy (anonymous):

oh , sorry my fault . the question is:

OpenStudy (anonymous):

What cubic model best fits this data?

OpenStudy (anonymous):

goshhh , i hate questions like these ! lol

OpenStudy (anonymous):

o wait, cubic models is something different... there is like a formula and stuff. Im not super familiar.

OpenStudy (anonymous):

yeah , me either really . . . . lol

OpenStudy (anonymous):

here, I looked it up and its long so give me a sec

OpenStudy (anonymous):

ok (:

OpenStudy (anonymous):

if you put you anwsers into ordered pairs, (5.027, 1908) ; (8.768, 1923) ; (12.29, 1960) ; (15.485, 1981). You can drop them into this system of equations. ax³ + bx² + cx + d = y for each pair. Got it?

OpenStudy (anonymous):

so you will have a(5.027)³ + b(5.027)² + c(5.027) + d = 1908 a(8.768)³ + b(8.768)² + c(8.768) + d = 1923 a(12.29)³ + b(12.29)² + c(12.29) + d = 1960 a(15.485)³ + b(15.485)² + c(15.485) + d = 1981

OpenStudy (anonymous):

shew.. then you just solve for each variable. which can be easily done with a ti-83.Or done painstakingly hard on paper.

OpenStudy (anonymous):

what about a ti-30 ?? lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ima plug it into my calc. If you feel like working that matrix then let me know

OpenStudy (anonymous):

still typing

OpenStudy (anonymous):

ok it is the last one